# Ina K. Schieferdecker Technische Universität Berlin, Einsteinufer 25, 10587 Berlin, Germany

### v2

@@ Abstract
The effectiveness of model-driven software engineering (MDSE) has been successfully demonstrated in the context of complex software; however, it has not been widely adopted due to the requisite efforts associated with model development and maintenance, as well as the specific modelling competencies required for MDSE. Concurrently, artificial intelligence (AI) methods, particularly deep learning methods, have demonstrated considerable abilities when applied to the huge code bases accessible on open-source coding platforms. The so-called big code provides the basis for significant advances in empirical software engineering, as well as in the automation of coding processes and improvements in software quality with the use of AI. The objective of this paper is to facilitate a synthesis between these two significant domains of software engineering (SE), namely models and AI in SE. The paper provides an overview of the current state of AI-augmented software engineering and develops a corresponding taxonomy, AI4SE. In light of the aforementioned considerations, a vision of AI-assisted Big Models in SE is put forth, with the aim of capitalising on the advantages inherent to both approaches in the context of software development. Finally, the new paradigm of pair modelling in MDSE is proposed.
@@ 1 Introduction
"Programming at the beginning of 2025 and at the end of 2025 are going to be completely different things."
Sam Altman, Febr. 7, 2025, The Age of AI – Panel Discussion, TU Berlin, Germany
@@
The advent of artificial intelligence (AI), and more specifically, generative AI, has precipitated a substantial transformation in the field of software engineering (SE), which is expected to undergo further significant changes in the near future (see also the quotation by Sam Altman, OpenAI). In order to better understand and comprehend these developments, this paper analysis the state of art in using AI for SE Software engineering is a field of informatics/computer science that addresses the development and analysis of systematic approaches for the design, development, verification & validation and maintenance of software and of software-based systems [^1] In this paper, the term ”software” will be used as a general reference to the collective term for computer programs and related SE artefacts.. It establishes principles and identifies optimal practices for the production and operation of software, with the aim of ensuring the reliability, security, scalability, maintainability and alignment of software with user, business and societal requirements. The objective of software engineering is the production of high-quality software that is cost-effective, delivered in a timely manner and can be readily adapted as requirements evolve.
@@
The term software engineering was coined in 1968 with the very first software crisis [82], but still software engineering practices and the resulting software quality have not kept pace with the quality levels required in critical domains or application contexts. In 2024, another dramatic story was added to the software horror show: The millions of enterprise outages caused worldwide by the CrowdStrike Falcon sensor update [24] were due to simple errors such as not validating the number of fields in an IPC template type, missing a runtime array bounds check, a logic error in the content validator, not testing template types enough, and not testing template instances within the content interpreter. This was unprofessional software development that ignored state of the art of SE.
However, software outages are not the only indication that SE still has no full answers to the challenges posed by the increasing complexity of software-based systems and the diversity of requirements placed on them: [71] analysed that "[t]here exists a statistically significant medium sized difference between open and closed source projects: the former have a DD [(defect density)] that is 4 defects per KLoC lower than the latter. Java projects exhibit a significantly lower DD than C projects, 4.1 defects per KLoC on average: In general the [code] size appears to be negatively correlated to DD: the larger the project the lower the DD. In particular, large projects are 10 times less defective than medium ones.". According to [59], the defect density (DD) of software projects within the industrial sector is estimated to fall within the range of 1 to 25. This suggests that a software program comprising one million lines of code (MLoC) may contain between 1,000 and 25,000 defects. Given the lines of code in a piece of software, deriving estimates of the approximate number of defects in it is a relatively straightforward process. To illustrate, [74] compares typical software such as operating systems, browsers, and office suites: The Windows 10 operating system has approximately 80 MLoC, Ubuntu 50 MLoC, MacOS X 84 MLoC, Android 12 MLoC, iOS 12 MLoC, the browser Google Chrome 6.7 MLoC, Mozilla Firefox 21 MLoC, or the office suites Microsoft Office 2013 45 MLoC, Apache OpenOffice 19 MLoC, and LibreOffice 10 MLoC. Therefore, the challenges associated with the development of accurate and suitable software remain significant.
@@
[17] argued that the primary challenge in software development lies in the specification, design, and testing of complex systems, rather than in the labour for representation of the system or testing its fidelity. It emphasized the distinction between essential difficulties (inherent complexity) and accidental difficulties (extraneous challenges), stating that past advances, like high-level programming languages, have only reduced the latter. The paper suggested that addressing accidental difficulties would not lead to major improvements in software development, as essential difficulties remain fundamental. [36] critiqued the notion of "accidental difficulties", arguing that these so-called accidents are often the result of negligence or poor practices, not mere happenstance. This paper advocated for a disciplined, science-based, and model-driven approach to software development, similar to traditional engineering disciplines. Building on this, [68] highlighted the growing societal dependence on autonomous and intelligent software systems, suggesting that new approaches are needed to ensure not only traditional software quality facets (safety, security, etc.), but also to address socio-technical and socio-political implications, particularly in human-machine collaboration.
@@
In model-driven development (MDD [70]), also known as model-driven software engineering (MDSE [69]), models are original artefacts that are engineered with the intention of facilitating the top-down construction of complex software. Furthermore, models are utilised during runtime to enable the monitoring, verification & validation of software operations, as well as the optimisation of its performance [41, 34, 85]. The issues in using traditional (top-down) MDSE has been e.g. analysed in [35].
As has been the case with the comprehensive incorporation of modelling in software engineering, AI techniques have been employed in software engineering activities from the outset, see for example [11]. In the context of significant advancements in AI-driven software engineering in 2021, Gartner’s hype cycle for emerging technologies [19] projected that AI-enhanced software engineering would reach the peak of inflated expectations within a five-to-ten-year timeframe, subsequently entering a phase of productivity stabilization. Since 2023, it is estimated that AI-augmented software engineering will reach the productivity plateau already in two to five years, while in previous iterations of the hype cycle, AI-augmented development and AI-assisted design were identified as emerging technologies on the rise. It was interesting to observe the evolving business expectations and the convergence of trends for AI-augmented design and development into a unified technological trend for AI-augmented engineering. Moreover, it is encouraging to observe that software engineering is receiving particular attention and prioritisation, which is indicative of the growing importance of this field.
@@
Therefore, this paper focuses on the intersection of AI and SE, specifically MDSE, to improve software engineering practices. The paper aims to bridge the gap between the theoretical advantages of MDSE and its practical limitations by integrating AI techniques, and introduces a novel taxonomy, AI4SE, to classify AI applications within the software engineering domain, highlight open research questions, and direct future research. The paper uses a combination of literature review, analysis of existing software engineering practices and AI applications, and development of a novel taxonomy to achieve its objectives. The main contributions of the paper include:
@@
  • The new AI in Software Engineering Taxonomy AI4SE. This taxonomy categorises AI applications based on the purpose of using AI, the target of AI within software engineering (e.g. requirements, design, coding, testing, operations, process), the type of AI used, and the level of autonomy achieved by the AI.
@@
  • The novel Modelling Hypothesis, which posits that models have statistical properties that can be used to develop more effective software engineering tools.
@@
  • The concept of Big Models, analogous to big code, as a basis for advanced empirical MDSE. The paper suggests that these models can be generated top-down, or extracted bottom-up from code, and that AI methods can be applied to these big models.
@@
  • The new paradigm of pair modelling in MDSE, where a software engineer and an AI tool/agent collaborate in software development, similar to pair programming. The aim is to leverage the strengths of both human and AI for MDSE to improve software quality.
@@
The AI4SE taxonomy is validated by categorising highly cited and recent publications in AI-assisted software engineering. This classification shows the various ways AI is being applied across different activities in SE, such as requirements engineering, architecture design, coding, testing, and deployment.
@@
In section 2, the paper addresses the role of models in software engineering. The evolution of big code on coding platforms is discussed in section 3 and the utilisation of AI for coding and software engineering in section 4. The novel taxonomy of AI for SE (AI4SE) to facilitate a deeper understanding of this evolving field is presented also in this section. Section 5 examines the practice of contributing software models to coding platforms, which is increasingly being done together with software code. In line with the increasing prevalence of large-scale models, it examines recent developments in AI-enhanced modelling and model utilisation within the domain of software engineering and presents the Big Model concept and the pair modelling paradigm. In section 6, related work is reviewed. A summary concludes the paper.
@@ 2 Models in Software Engineering
According to the IEEE Software Engineering Body of Knowledge [73] (see Figure 1), the software lifecycle is comprised of distinct phases and activities that can be described as knowledge areas, including requirements, architecture, design, construction, testing and maintenance. There are also knowledge areas that deal with the fundamentals of computer science, mathematics and engineering, as well as cross-cutting activities pertaining to software quality, security, configuration management and engineering management. Furthermore, there are cross-cutting knowledge areas that encompass the processes, operations, professional practice and economics of software engineering, along with a distinct area for software engineering models and methods. This knowledge area encompasses the processes and methodologies associated with modelling, the various types of models, including those pertaining to information, behaviour and structure, and the analysis of models, see also [69]. Overall, SWEBOK addresses dedicated model-based or model-driven approaches in each knowledge area, such as model-based requirements specification, model-based architecture, model-driven design, lifecycle models, executable models, model-based testing, etc.
@@
This is because models in SE are the key to dealing with the complexity of software: Models provide the essential abstractions to capture requirements, support design decisions, or to offer comprehensive overviews of software structures and behaviours. They are essential tools in the engineering process for constructing and maintaining software. They are also vital for configuration, monitoring and other runtime support during software operation and management. The understanding of Model-Driven Software Engineering (MDSE) has evolved in recent years as the capabilities of computing, modelling and the available software and data have evolved:

### v3

@@ Abstract
The effectiveness of model-driven software engineering (MDSE) has been successfully demonstrated in the context of complex software; however, it has not been widely adopted due to the requisite efforts associated with model development and maintenance, as well as the specific modelling competencies required for MDSE. Concurrently, artificial intelligence (AI) methods, particularly deep learning methods, have demonstrated considerable abilities when applied to the huge code bases accessible on open-source coding platforms. The so-called big code provides the basis for significant advances in empirical software engineering, as well as in the automation of coding processes and improvements in software quality with the use of AI. The objective of this paper is to facilitate a synthesis between these two significant domains of software engineering (SE), namely models and AI in SE. The paper provides an overview of the current state of AI-augmented software engineering and develops a corresponding taxonomy, ai4se. In light of the aforementioned considerations, a vision of AI-assisted big models in SE is put forth, with the aim of capitalising on the advantages inherent to both approaches in the context of software development. Finally, the pair modelling paradigm is proposed for adoption by the MDSE industry.
@@ 1 Introduction
"In 12 months, we may be in a world where AI is writing essentially all of the code."
Dario Amodei, CEO Anthropic, Mar. 10, 2025
@@
The advent of artificial intelligence (AI), and more specifically, generative AI, has precipitated a substantial transformation in the field of software engineering (SE), which is expected to undergo further significant changes in the near future (see also the quotation by Dario Amodei, Anthropic). In order to better understand and comprehend these developments, this paper analyses the state of art in using AI for SE Software engineering is a field of computer science that addresses the development and analysis of systematic approaches for the design, development, verification, validation, and maintenance of software and of software-based systems [^1] In this paper, the term ”software” will be used as a general reference to the collective term for computer programs and related SE artefacts.. It establishes principles and identifies optimal practices for the production and operation of software, with the aim of ensuring the reliability, security, scalability, maintainability and alignment of software with user, business and societal requirements. The objective of software engineering is the production of high-quality software that is cost-effective, delivered in a timely manner and can be readily adapted as requirements evolve.
@@
The term software engineering was coined in 1968 with the very first software crisis [114], but still software engineering practices and the resulting software quality have not kept pace with the quality levels required in critical domains or application contexts. In 2024, another dramatic story was added to the software horror show: The millions of enterprise outages caused worldwide by the CrowdStrike Falcon sensor update [33] were due to errors such as failing to validate the number of fields in an IPC template type, omitting a runtime array bounds check, a logic error in the content validator and insufficient testing of template types and template instances within the content interpreter. Such software development, which ignores the state of the art in SE, can be described as deficient, if not unprofessional. Software failures are only one indication that SE still lacks comprehensive solutions to the challenges posed by the increasing complexity of software-based systems and the diverse requirements placed on them. The sheer number of software defects, as discussed in [98], also serves as an indicator. This raises the question of whether AI can not only be a very useful tool for SE, but can also contribute to promoting the latest SE research results, for example for MDSE.
@@
This paper therefore discusses the current state of AI applications in SE, how various approaches can be categorised and connected, and how AI could help implement MDSE to further improve software quality: Already [25] argued that the greatest challenge in software development lies in the specification, design and testing of complex systems, rather than in the work involved in representing the system as code. It emphasized the distinction between essential difficulties (inherent complexity) and accidental difficulties (extraneous challenges), stating that past advances, like high-level programming languages, have only reduced the latter. The paper suggested that addressing accidental difficulties would not lead to major improvements in software development, as essential difficulties remain fundamental. [47] critiqued the notion of "accidental difficulties", arguing that these so-called accidents are often the result of negligence or poor practices, not mere happen stance. This paper advocated for a disciplined, science-based, and model-driven approach to software development, similar to traditional engineering disciplines.
@@
In model-driven development (MDD [97]), also known as model-driven software engineering (MDSE [95]), models are original artefacts that are engineered with the intention of facilitating the top-down construction of complex software. Furthermore, models are utilised during runtime to enable the monitoring, verification, and validation of software operations, as well as the optimisation of its performance [52, 45, 119]. The core idea behind using models in SE is to exploit their abstraction and comprehensibility. In theory, models should be easier to develop, use, document and maintain. If code could be fully generated from models automatically, work could be done solely at the modelling level. However, several studies have analysed issues with MDSE, including the integration of generated and hand-written code, the modularity of the generated code, and the organisation of the development process (see [46], for example).
In another line of research and development, AI techniques have been employed in software engineering activities from the outset, see for example [19]. Against the backdrop of significant advances in AI-assisted software development, [26] predicted that it would reach the peak of inflated expectations within five to ten years, after which it would enter a phase of productivity stabilisation. Subsequently, [29] estimated that AI-assisted software development would reach the productivity plateau within just two to five years. The ongoing increase in demand for software, driven by the global shift towards digitalisation, can be met by utilising AI solutions in software engineering. Despite the support of AI, software engineering continues to be important for the development of high-quality, fit-for-purpose software. The focus is on augmentation, not replacement, of software engineers.
@@
Therefore, this paper investigates the intersection of AI and SE, specifically MDSE, to improve software engineering practices. The paper aims to bridge the gap between the theoretical advantages of MDSE and its practical limitations by integrating AI techniques. It introduces a novel taxonomy, ai4se, which classifies research on AI methods within the SE domain, as well as the more specific MDSE domain. The paper draws on a literature review and an analysis of SE practices and AI applications to develop the taxonomy. The main contributions of the paper include:
@@
  • The new AI in Software Engineering Taxonomy ai4se. This taxonomy categorises AI applications based on the purpose of applying AI, the target of AI within SE such as requirements engineering or the testing process, the type of AI used, and the level of autonomy achieved by the AI.
@@
  • The new Model Naturalness Hypothesis, which posits that models have statistical properties that can be used to develop more effective software engineering tools.
@@
  • The new concept of Big Models, analogous to big code, as a basis for advanced empirical MDSE. The paper suggests that these models can be generated top-down, or extracted bottom-up from code, and that AI methods can be applied to these big models.
@@
  • The renewed paradigm of pair modelling in AI-augmented MDSE involves a software engineer and an AI tool/agent collaborating on software development in a manner similar to pair programming. The aim is to combine the strengths of humans and AI to improve the quality of software and enable the immediate consideration of ethical issues in MDSE.
@@
The ai4se taxonomy is validated by categorising highly cited and recent publications in AI-augmented SE. This classification shows the various ways AI is being applied across different activities in SE, such as requirements engineering, architecture design, coding, testing, and deployment.
@@
Section 2 discusses the availability of large code sets on coding platforms that form the basis for training AI using actively applied, operated and maintained code. Section 3 addresses the role of models in SE. Bringing these two sections together, Section 4 discusses the utilisation of AI for SE. This section presents the taxonomy of AI for SE ai4se, which has been developed to facilitate a deeper understanding of this evolving field. Finally, Section 5 discusses the adoption of MDSE practices with the help of AI. This section examines recent developments in AI-augmented modelling and model utilisation within SE, presenting the big model concept and the pair modelling paradigm in AI-augmented MDSE. Section 4 reviews related work. The paper concludes with a summary and outlook.
2 Big Code in Software Engineering
The online availability of open-source software (OSS) paved the way for software reuse. It has also transformed the landscape of software licences and the way software is developed, with developers using OSS for specific components or entire products. OSS has also enabled deeper insights into software architecture and development practices. It has enabled new approaches to empirical software engineering, deepening our understanding of SE.
The advent of software coding platforms can be traced back to the 1990s, although it was not until the early 2000s that they truly gained prominence, with the launch of GitHub in April 2008 representing a significant milestone in this regard. At the present time, GitHub is the most widely utilised software coding platform. As stated by [23], GitHub hosts over 100 million projects and 40 million users. It became not only a significant platform for software engineering collaboration, but also a prominent reference for open-source software mining [58]. The study also demonstrated that a considerable number of GitHub repositories are not directly related to software development. This is because GitHub is not solely utilised for coding collaborations; it is also employed for collaborations on websites, editing books or other publications, and is even used as a storage platform. A later manual analysis by the same authors [59] revealed that over a third of the repositories on GitHub were not software development repositories. According to [10], code repositories constitute around 20% of the lines stored on GitHub (see Figure 2(a)).
As the vast quantities of software code have accumulated over time, the term "Big Code" has evolved (see, for example, [78, 86, 116]): Not only can the available code be reused, extended and analysed, it can also be used to train ML models, enabling the application of AI for SE on a large scale [69]. The term big code signals a fundamental shift in the analysis and construction of software, moving from a focus on the formal logic of a single program to identifying statistical patterns within vast repositories of source code. [7] considered the statistical properties of code to be similar to those of natural language, meaning code is full of learnable, predictable patterns. This observation later led to the formulation of the naturalness hypothesis [8].
@@ 3 Models in Software Engineering
According to the IEEE Software Engineering Body of Knowledge [100] (see Figure 1), the software lifecycle is comprised of distinct phases and activities that can be described as knowledge areas, including requirements, architecture, design, construction, testing, and maintenance. There are also knowledge areas that deal with the fundamentals of computer science, mathematics, and engineering, as well as cross-cutting activities pertaining to software quality, security, configuration management, and engineering management. Furthermore, there are cross-cutting knowledge areas that encompass the processes, operations, professional practice, and economics of software engineering, along with a distinct area for software engineering models and methods. This knowledge area encompasses the processes and methodologies associated with modelling, the various types of models, including those pertaining to information, behaviour, and structure, and the analysis of models, see also [95]. Overall, SWEBOK addresses dedicated model-based and model-driven approaches in each knowledge area, such as model-based requirements specification, model-based architecture, model-driven design, lifecycle models, executable models, model-based testing, etc.
@@
This is because models in SE are the key to dealing with the complexity of software: Models provide the essential abstractions to capture requirements, support design decisions, or to offer comprehensive overviews of software structures and behaviours. They are essential tools in the engineering process for constructing and maintaining software. They are also vital for configuration, monitoring, and runtime support during software operation and management.
However, the adoption of MDSE in the industrial practice remains limited. [6] reports that, in this survey, only 37.4% of the participants stated that "MD[S]E is used in most projects and by most developer", while 46% reported limited or no use. Nowadays, however, MDSE does not only rely on models in the Unified Modelling Language (UML), in other standardized modelling languages, such as Business Process Modelling Language (BPMN), or in Domain-Specific Languages (DSL): Rather than explicitly developing software models and using them from the top down, they can be extracted from software executions, monitoring, or static analysis. These approaches aim to minimise the costs associated with model development while still reaping the benefits of MDSE [14]. The combination of explicitly developing model artefacts and extracting models from code artefacts has resulted in a hybrid top-down/bottom-up MDSE approach, as described in [104]. This approach benefits from both directions of MDSE, as demonstrated in [115, 119, 48]. Furthermore, as large volumes of data are interpreted more dynamically depending on their current application context, new data modelling approaches have evolved [90]. Examples include Not Only SQL (NoSQL), JavaScript Object Notation (JSON), Resource Description Framework (RDF), and Yet Another Markup Language (YAML).
The formal structures of models, together with the increased use of advanced data models and the bottom-up and hybrid modelling approaches, have paved the way for MDSE with AI by making larger model sets available for AI training. Additionally, machine learning is employed to render models suitable for further processing, even when they are not inherently machine-processable, as with images: [99] presents an approach to automatically detecting images containing one or more of the ten types of UML diagrams being analysed. This approach was used to derive a larger model set containing these ten types of UML diagrams. It complements other approaches for providing ground truth model sets, such as the Lindholmen model set [91, 92], which includes ’almost 100k models from 22k [GitHub] projects’, and the ModelSet [75].
(a) Lines and portion of GitHub code repositories
(b) Number and percentage of GitHub files in code repositories
(c) Number of modelling files in code repositories
An analysis using the GitHub search API indicates that around 1% of GitHub files contain models or partial model information. Counting the number of code and model files in GitHub repositories reveals the ratio of code to model files for software designs, interface models, and data models (see Figure 2(b)). Figure 2(c) shows the split between design, interface and data models. These figures are only rough indicators. Lines of code and the number of files are crude measures of software. File extensions do not accurately indicate file types. Furthermore, there are numerous coding and modelling tools, some of which have proprietary file extensions. Images are not counted.
Nevertheless, the analysis showing a low number of modelling files on GitHub is consistent with the analysis of [46] from over a decade ago. The authors of this study argue that, while code generators can produce large amounts of code, not all of it is necessarily useful or would have been written manually. The effort required to develop, maintain and adapt the generators offsets the productivity gains from modelling. Manual code changes make round-trip engineering difficult. Generated code must be checked by developers to ensure correctness, which limits the benefits of automation. Also, poor software design models can lead to inefficient development. Consequently, developers often prefer generic solutions with hand-written, reusable code to customised, generated solutions. Furthermore, as programming languages and frameworks become more expressive, the need for model-driven code generation would diminish. However, while generative AI can empower software developers further, it also makes MDSE more practical and applicable to a wider range of software projects. The ability to regenerate models from code and/or keep them in sync with evolving software requirements and manual code changes addresses the limitations of MDSE that have been identified thus far. Despite these limitations, the Lindholmen dataset and the GitHub analysis demonstrate that MDSE has been used continuously for the aforementioned reasons. However, the GitHub analysis also reveals that modelling is currently predominantly employed for data and interface structures, as robust functionality of these elements hinges on a well-defined structure (see, for instance, [33]). The analysis also shows that design and behaviour models for software projects are rarely used.
4 Artificial Intelligence in Software Engineering
Big code (see Section 2) constitutes the basis to train AI in recognising coding patterns and anti-patterns. These are now accessible via AI tools that can complete code, detect defects, summarise code, translate code, enhance code searches, and check and learn coding conventions, among other things. The potential of using AI for SE was discussed as early as 1988 in [19], and has been a topic of discussion ever since, most recently for instance in [5]. For example, [43] proposes the AI in SE Application Levels (AI-SEAL) taxonomy, which differentiates between the point of applying AI to the software engineering process, the software product or at runtime, the levels of automation, and the types of AI along the five tribes differentiation by [36]. It is noteworthy that the majority of the papers analysed in [43] employ AI in the software engineering process. However, despite the numerous facets of the software engineering process (see Figure 1), this study does not provide further elaboration. Nevertheless, other publications offer more detailed discussions of AI in SE, including [18] or [87], see also Section 6. Notwithstanding the aforementioned considerations, a taxonomy for the role of AI in software engineering has yet to emerge.
In light of the recent advancements in machine learning (ML), in particular in deep learning (DL), that have made significant breakthroughs possible, this paper focuses mainly on the latest developments in the application of ML to SE. For ML to be effective, it is essential to utilise the appropriate structures inherent to SE artefacts, which has been a topic of considerable debate: [8] presents an overview of the various ways in which source code can be represented, including representational models of tokens, token contexts, program dependency graphs, API calls, abstract syntax trees, object usage, and others. These representational models are used for AI assistance in SE, including the creation of recommender systems, the inference of coding conventions, the detection of anomalies and defects, the analysis of code, the rewriting and translation of code, the conversion of code to text for the purposes of documentation and information retrieval, and the synthesis and general generation of code from text. [61] adds further applications of AI assistance such as code completion, API migration and code repair. Furthermore, [8] presents "[t]he naturalness hypothesis[:] Software is a form of human communication; software corpora have similar statistical properties to natural language corpora; and these properties can be exploited to build better software engineering tools.". Given the intrinsic formats and formal characteristics of coding and modelling languages employed in SE, the second aspect of the naturalness hypothesis can be considered relatively straightforward. Yet, the initial proposition reinforces the necessity for SE artefacts that are readily comprehensible. This assertion was previously made by [44] in a different form: "[A]ny fool can write code that a computer can understand, good programmers write code that humans can understand.". Consequently, the application of AI in SE is not merely concerned with code generation or code repair; it also encompasses the enhancement of code through techniques such as refactoring, with the objective of optimising readability and maintainability.
The purpose dimension has the three facets of (P1) understanding to gain insight, (P2) generating to derive (parts of), and (P3) improving to refine SE artefacts and/or SE processing. An approach may address multiple purposes. If it is hybrid in this sense, all its purposes are made explicit, e.g. to support understanding only is denoted by ’P1’, or to support both understanding and generation is denoted by ’P1, P2’.
The ai4se application targets consist of the main SE activities as described by SWEBOK [100] and DevOps [65]: (S1) Software development includes requirements engineering, architecture and design, coding, testing, and maintenance. (S2) Software operations includes deployment, operation, and monitoring. Together with the (S3) SE process including management and improvement, the three form the target of application dimension in the taxonomy. Each of the target options can be performed with the support of models, based on models or even driven by models, or without them. For the sake of comprehension of the taxonomy, the model option is optional for each application target: If models are used, i.e. the application target is MDSE, it is marked with "(m)". Thus, the application target "S1.m" refers to software development with models, and "S1" to one without. Or, in a more fine-grained analysis, "S1.d.m" would be used for model-based testing and "S1.d" for testing without models. An approach may address several application targets. If it is hybrid in this sense, all of its goals can be made explicit, e.g. support for coding only is denoted by "S1.c", support for both coding and testing is denoted by "S1.c, S1.d", or support for coding and model-based testing is denoted by "S1.c, S1.d.m". For simplicity, "h" can also be used to indicate a hybrid approach of several targets.
The types of AI are divided into (T1) symbolic AI, i.e. knowledge representation, rule-based systems, and logical reasoning, (T2) being subsymbolic including statistical AI, i.e. probabilistic modelling and inference, Bayesian networks, Markov models, Gaussian processes, etc., classical machine learning, i.e. supervised, unsupervised, and reinforcement learning, feature engineering and selection, etc., and deep learning, i.e. neural networks, transformers, transfer learning, self-supervised learning, attention mechanisms, etc., (T3) generative AI (GenAI), i.e. large language models (LLM), generative adversarial networks, diffusion models, variational autoencoders, autoregressive models, contrastive learning, reinforcement learning from human feedback, etc., (T4) agentic AI, i.e. multi-agent systems, planning and decision making, hierarchical reinforcement learning, etc., and (T5) general purpose AI, i.e. meta-learning and world models. Today, all types (T1) to (T5) are used in software engineering, but not (T5) due to its non-existence. An approach can also use a combination of AI types, i.e. it can be hybrid. This can be made explicit, e.g. by denoting "T1, T2" for a hybrid approach that uses both symbolic and classical machine learning. For simplicity, "h" can just be used to denote a hybrid approach that uses more than one type of AI.
The levels of autonomy/automation are defined similarly to the levels of autonomous driving [17] as five levels from no support to recommendations to full automation: (L0) No AI-assisted automation, (L1) AI-assisted options, where the developer is in full control and receives recommendations to choose from and adapt, (L2) AI-assisted selection, where the AI selects pre-defined options, (L3) AI-based partial automation, where the AI selects options in simple, standard cases, and (L4) AI-based full automation, where the AI operates without the developer. So far, level L1 and L2 are the most common, level L3 is on the rise and level L4 is rather unrealistic for complex, industrial-scale software. Higher levels of autonomy typically include the capabilities of lower levels of autonomy; when presenting the level of autonomy of an approach, the highest level of autonomy is given.
(a) Overview of the ai4se ontology
(b) Sample paper classified in ai4se
(c) Target dimension of the ai4se ontology
Figure 3: The ai4se taxononmy defined as an ontology
To define this taxonomy more formally, the Web Ontology Language (OWL) [79] was used, due to its association with the Semantic Web and the support offered by the Protégé tool [84] (see Figure 3). The ai4se ontology definition is based on gUFO [9] as a grounding and uses, in addition, the Computer Science Ontology (CSO [94]) to associate a general research topic with a research paper. A research paper itself is an endurant and the dimensions used to classify its research contribution are abstract individual types. However, as the ontology lends itself to interactive elaboration rather than graphical representation, this paper uses a different tabular representation to depict the classification of the research papers, as shown in Figure 4. Interested readers can, however, explore the classification of research papers directly in the ontology [96], to which they can also contribute. The ai4se ontology can be used not only to understand the concepts in the research field of AI for SE better, but also to explore research related to a specific aspect, such as all papers on agentic AI for SE, using SPARQL queries as shown in Listing 1 and 2222SPARQL is the standard query language and protocol for linked open data on the Web as defined by RDF triples. The OWL ontology ai4se is already encoded in RDF triples, making it immediately queryable with SPARQL.. ai4se has been defined as an ontology because, although taxonomies have been proposed for every area of SE knowledge within the SWEBOK [100] as discussed in [113], neither SWEBOK nor any other source defines a comprehensive SE ontology that could be used to classify research results for the purpose of this paper. For example, while the CSO is quite detailed, it does not cover the knowledge areas of SWEBOK. For example, it lacks the concept of ’runtime monitoring’ of software components and includes ’object-oriented design’ at the top level, even though this is just one type of software design.
1PREFIX rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#>
2PREFIX ai4se: <http://purl.org/ai4se/ontology#>
3SELECT paper
4WHERE{
5 paperai4se:hasLevel ai4se:AI-assisted_options .
6}
Listing 1: All papers on AI-assisted options automation.
1PREFIX rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#>
2PREFIX ai4se: <http://purl.org/ai4se/ontology#>
3SELECT DISTINCT target
4WHERE{
5 paperrdf:type ai4se:ResearchPaper .
6 paperai4se:hasTarget target.
7}
Listing 2: All software engineering targets addressed by research papers.
Figure 4: Overview of the ai4se taxonomy
Paper Summary Purpose Target Type Level
[71] T-BERT generates trace links between source code and natural language artefacts. It represents a significant advancement in automated software traceability, offering improved accuracy, efficiency, and practicality. P2 S1 T2.c L3
[67] Natural language processing (NLP) approach for improving requirements engineering in the automotive industry: It identifies flaws, such as ambiguous words, incomplete process descriptions, or incorrect quantifiers. P3 S1.a T1 L2
[88] CBRank method for prioritising software requirements: It uses stakeholder input and requirement attributes to generate an optimal ranking and enables adaptive elicitation by refining the prioritisation based on real-time feedback. P3 S1.a, S2.a, S3.a T1, T2.b L2
[20] ADeX is a tool designed to automate the curation of architectural design decisions in software development. It helps automate the management of architectural decisions, reducing documentation effort while improving decision accuracy. P1, P3 S1.b, S2.a T1, T2.b L3
[22] Explores the early experiences of developers using the AI-augmented coding assistant GitHub Copilot. It may expand to debugging, code review, and software maintenance, requiring better trust and provenance tracking. P1, P2 S1.c T3 L3
[106] IntelliCode Compose is a code completion tool that uses a generative transformer model to predict sequences of code tokens, including entire lines of code. It has been trained on a massive dataset of 1.2 billion lines of code across multiple programming languages. P2, P3 S1.c T2.a, T2.c, T3 L3
[13] Getafix is an automated bug-fixing tool that learns from past, human-written fixes to suggest human-like fixes for static analysis warnings. P2, P3 S1.c, S1.d T2.a, T2.b L2
[49] GraphCodeBERT improves code understanding by incorporating dataflow graphs into the BERT transformer architecture. Data flow graphs represent dependencies between code variables and improve semantic understanding over treating code as a simple sequence of tokens. P1 S1.c, S1.d, S1.e T2 L1
[51] DeepCodeReviewer is an automatic code analysis system that learns from historical peer reviews to identify and suggest relevant reviews for code snippets. P1, P3 S1.c, S1.d T2 L2
[37] Introduces DeepDebug or automated bug detection and fixing in Java: It generates non-deletion fixes that modify instead of just deleting code P1, P2 S1.c, S1.e T3 L3
[112] Approach improves code review automation with DL by working with pre-trained models on raw source code and English text. P1 S1.c, S1.d T3 L2
[108] The book discusses the application of AI beside others for infrastructure monitoring and anomaly detection, predictive maintenance, and build pipeline optimization. P1, P3 S1.c, S1.d, S2.a, S3.b T2.a, T2.b L3
[24] Presents RepairAgent for automated program repair that allows to autonomously plan and execute actions, such as gathering information, searching for code snippets, and validating fixes, by interacting with a set of specifically designed tools. P3 S1.c, S2.a T4 L3
[68] Approach links different types of testing information to help automate the generation of equivalence classes for testing. The linking supports tasks such as test reduction, prioritisation and selection. P1, P2, P3 S1.d T2.a, T2.b L2
[125] Presents a hybrid approach to test case prioritization for regression testing aiming to minimise testing costs. Aims to enhance software testing efficiency by optimising test case execution order. P3 S1.d, S1.e T1, T2.a, T2.b L2
[102] Introduces Retecs for automatic test case prioritisation and selection in continuous integration. It uses test case meta-data to adapt to changes in the environment and prioritise error-prone test cases. P3 S1.d, S2.a T2.b, T2.c L3
[28] Explores machine learning-based ensemble methods to predict software quality more accurately than previous approaches P1 S1 T2.b L3
The fidelity of the new taxonomy ai4se (Figure 4) is verified by reviewing and classifying highly cited publications, but also recent publications to demonstrate the applicability of this taxonomy. A lightweight systematic literature review (SLR), as described in [103], was conducted to analyse related research. This SLR protocol was followed:
  • Review title: SLR on evaluating the status of AI for SE research by applying the ai4se taxonomy.
  • Objectives of the review: